# Foundation Models for Remote Sensing and Earth Observation: A Survey

### v2

@@ 6.2 Generative Models for RS
However, the unique characteristics of RS data—often multi-spectral, irregularly sampled over time, and involving complex spatio-temporal dependencies—require specialized approaches that go beyond generative FMs developed for natural images. Despite these challenges, generative models hold great promise for RS applications such as super-resolution, cloud removal, and temporal inpainting, underscoring the need for tailored generative FMs in the RS domain.
Recently, Khanna et al. [137] introduced DiffusionSat, the first generative FM specifically designed for satellite imagery, inspired by the architecture of Stable Diffusion. DiffusionSat uses metadata commonly associated with satellite imagery—such as latitude, longitude, timestamps, and GSD —along with textual descriptions to train the model for single-image generation from publicly available satellite datasets. Additionally, the model includes conditioning mechanisms that allow for fine-tuning across specific generative tasks and inverse problems, such as super-resolution, inpainting, and temporal generation. Additionally, Zheng et al. [138] develop generative probabilistic change model based on diffusion model for change detection.
6.3 Weather Forecasting Foundation Models
6.4 Summary and Discussion
7 Benchmark Performance
In this section, we compare, analyze, and discuss the existing RSFM studies reviewed in this survey. It is important to note that these studies have distinct research focuses, modalities, learning setups, and target tasks or applications, making it difficult to conduct fair comparisons using unified benchmarks. Furthermore, as RSFM development is still in its early stages, robust and comprehensive evaluation frameworks across different tasks and modalities are yet to be fully established. Given these limitations, we focus on key dimensions and try to provide an informative analysis and insights into the current progress and future directions.
7.1 Performance of VFM Pre-training
As outlined in Section 4.1, VFM pre-training models are subsequently fine-tuned for various downstream RS tasks [140]. Table IV presents evaluations for seven widely adopted downstream tasks across different modalities, including: BigEarthNet [11] for multi-label land cover classification; EuroSAT [141] for LULC classification; Onera Satellite [142] for change detection tasks; SpaceNet [143] for building segmentation; Functional Map of the World (FMoW) [10] for high-resolution satellite image time series classification among 62 categories; DIOR-H [144] for horizontal object detection and DIOR-R [145] for oriented object detection. All metrics are taken from the respective papers. We select the most widely used datasets for this comparison, though other benchmarks, such as Geo-Bench [146], along with those tailored for different tasks, could also provide valuable insights.
@@ 7.1 Performance of VFM Pre-training
7.2 Performance of VLM Transfer Learning
This subsection compares the zero-shot performance of various VLMs in RS. Although the field remains relatively underexplored, we source benchmark results from [116] as in Table V, which evaluated different types of FMs, including general-domain VLMs like CLIP and RS-specific models fine-tuned from CLIP, such as CLIP-RSID22footnotemark: 2, RemoteCLIP[105], and GRAFT[116], across typical RS tasks like image classification and retrieval.
The findings show that while CLIP demonstrates some effectiveness on RS datasets, its performance falls short compared to its success on natural image benchmarks. In contrast, recent RS-specific VLMs exhibit significant improvements, underscoring the potential of tailoring VLMs for RS applications. However, these specialized models tend to perform well on specific benchmarks but show less overall robustness compared to CLIP. These results highlight both the progress made and the substantial opportunities for future advancements in this domain.
8 Future Directions
@@
Larger and More Diverse Datasets: FMs thrive on vast amounts of data, with the principle being ”the more data, the better” [1]. Building large, diverse datasets for RSFMs is foundational, but annotating multimodal RS data poses significant challenges. As the volume of RS data grows, effective data management—including integration, privacy, governance, and quality control—becomes increasingly important. While progress has been rapid, we foresee a need for even larger and more comprehensive datasets to drive the next phase of RSFM development.
Multimodal RSFM: An ideal RS interpretation system should be capable of processing multiple sensor modalities simultaneously, providing a richer and more holistic view of geospatial data. Beyond sensor modalities, RSFMs should also incorporate other data types—such as text, dense annotations for segmentation, and bounding boxes for object detection—enabling them to handle a broad spectrum of tasks with greater flexibility.
@@
High Processing Speed: In time-sensitive EO applications such as disaster monitoring, real-time data processing is crucial for effective response and decision-making. RSFMs shall be designed to handle vast amounts of RS data with high efficiency, enabling rapid analysis and forecasting. This will ensure that RSFMs can provide actionable insights swiftly, supporting critical tasks like early warning systems, damage assessment, and emergency response coordination.

### v3

@@ VII Benchmark Performance
VII-A Performance of VFM Pre-training
As outlined in Section IV-A, VFM pre-training models are subsequently fine-tuned for various downstream RS tasks [157]. Table V presents evaluations for seven widely adopted downstream tasks across different modalities, including: BigEarthNet [14] for multi-label land cover classification; EuroSAT [158] for LULC classification; Onera Satellite [159] for change detection tasks; SpaceNet [160] for building segmentation; Functional Map of the World (FMoW) [13] for high-resolution satellite image time series classification among 62 categories; DIOR-H [161] for horizontal object detection and DIOR-R [162] for oriented object detection. All metrics are sourced from the respective papers. We select the most widely used datasets for this comparison, though other benchmarks, such as Geo-Bench [163], along with those tailored for different tasks, could also provide valuable insights.
@@ VII-A Performance of VFM Pre-training
VII-B Performance of VLM Transfer Learning
This subsection evaluates the zero-shot performance of various VLMs in RS. We begin with a closed-set evaluation, comparing general-domain VLMs like CLIP with RS-specific models fine-tuned from CLIP, such as CLIP-RSID22footnotemark: 2, RemoteCLIP [96], and GRAFT [124], on key RS tasks like image classification and retrieval. These models enable zero-shot evaluation by projecting images and text into a shared embedding space. The process involves extracting image features via an image encoder and generating text embeddings from class-specific prompts (e.g., arbitrary category names) using a text encoder. The model then calculates cosine similarity between the image and text embeddings, assigning the label with the highest similarity score. Performance is measured using standard classification metrics, allowing models to generalize across diverse tasks without additional task-specific training.
Table VI present benchmark results sourced from [124]. It shows that while CLIP demonstrates some effectiveness on RS datasets, its performance falls short compared to its success on natural image benchmarks. In contrast, recent RS-specific VLMs exhibit significant improvements, underscoring the potential of tailoring VLMs for RS applications. However, these specialized models tend to perform well on specific benchmarks but show less overall robustness compared to CLIP. These results highlight both the progress made and the substantial opportunities for future advancements in this domain.
Another type of benchmarks base on VQA for open-set evaluation. These models perform zero-shot evaluation by encoding images and questions separately, then fusing their representations to understand the relationship between visual and textual information. Based on this joint representation, different model types either generates an answer in an open-ended format or selects the most probable answer from a predefined set.
As an example of open-set benchmark, GEOBench-VLM [164] was recently introduced for evaluating VLMs in RS, encompassing 31 fine-grained tasks across eight broad categories, including scene and object classification, object detection, segmentation, captioning, event detection, and non-optical and temporal understanding tasks. This benchmark provides a comprehensive assessment of perception tasks for VLMs in geospatial applications. Figure 11, presents evaluations of 10 open-source and advanced proprietary VLMs. The results reveal a significant performance gap between open-source and closed-source models, while all models exhibit suboptimal performance across various geospatial recognition tasks. The primary limitation stems from data biases in training these models, as current large-scale multimodal datasets contain limited geospatial data, restricting their ability to generalize effectively to RS-specific tasks. This highlights the need for more advanced and specialized models to support complex Earth observation and geospatial interpretation tasks, underscoring the importance of dedicated geoscience and remote sensing efforts in building more powerful and domain-adapted RSFMs to bridge this gap.
VIII Future Directions
@@
Larger and More Diverse Datasets: FMs thrive on vast amounts of data, with the principle being ”the more data, the better” [1]. However, the current RS datasets fall short of supporting the training of truly large-scale RSFMs. A notable observation is the absence of the “emergent capabilities” observed in general-domain FMs, indicating that the existing scale and diversity of RS data are insufficient to enable comparable advances. In parallel, although RSFMs fine-tuned from general FMs often yield superior performance on specific benchmarks, they still fall short in generalizing to out-of-distribution data.
Future dataset construction efforts should not focus solely on increasing data volume, but also prioritize critical dimensions such as modality diversity, spectral and temporal resolution, and dataset biases. As summarized in Tables II and IV, most available datasets are still centered on optical RGB and MSI. While there is some variation in GSD, the majority of studies rely on mid-resolution, publicly available sources like Sentinel imagery. To support more complex geospatial tasks, there is a pressing need for higher-resolution datasets with richer contextual information. Additionally, although recent initiatives have started to incorporate temporal sequences for RS change detection, significant progress is still needed.
Moreover, annotating large-scale, multimodal RS datasets remains a major bottleneck due to the inherent complexity and heterogeneity of the data. As the scale of RS data continues to grow, robust data management practices, including integration, quality assurance, governance, and privacy controls, will become increasingly critical. Despite rapid advancements, we anticipate that the next stage of RSFM development will require datasets that are not only larger and more diverse, but also more structurally and semantically enriched.
Multimodal RSFM: An ideal RS interpretation system should be capable of processing multiple sensor modalities simultaneously, providing a richer and more holistic view of geospatial data. Beyond sensor modalities, RSFMs should also incorporate other data types, such as text, dense annotations for segmentation, and bounding boxes for object detection, enabling them to handle a broad spectrum of tasks with greater flexibility.
@@
Scalability and Computational Efficiency: The scalability of RSFMs remains a key challenge due to the high computational cost associated with processing large, multi-modal, and high-resolution RS data. Compared to conventional vision tasks, RS applications often demand greater memory and processing power, particularly for multi-temporal and hyperspectral analysis. These challenges are amplified in scenarios requiring on-board processing on satellites, UAVs, or edge devices, where computational resources are limited. Despite their importance, few studies have explored model compression techniques such as pruning, quantization, or knowledge distillation, for tailoring large RSFMs to lightweight, efficient variants. Such strategies are critical to improving adaptability and deployment feasibility. In time-sensitive EO applications like disaster response and environmental monitoring, RSFMs must support real-time or near-real-time inference. Future work should prioritize optimizing RSFMs to balance performance with efficiency, enabling timely, scalable, and actionable geospatial intelligence.
@@
Reasoning for Geospatial Tasks: Most existing RSFMs primarily focus on perception-oriented tasks such as scene classification, object detection, and semantic segmentation, with only limited extensions to relatively simple reasoning functions (e.g., spatial relation judgment). These capabilities form the foundation of many geospatial applications. However, real-world EO tasks often require higher-order reasoning that goes beyond direct recognition. For example, analysts may need to integrate multi-temporal observations, infer causal relationships (e.g., attributing deforestation to specific policy or climate drivers), or correlate multimodal data sources to detect abnormal patterns or changes. These tasks demand spatial reasoning (e.g., understanding topological relationships), temporal reasoning (e.g., trend forecasting or event progression), and causal reasoning (e.g., identifying cause-effect dynamics in land use change). Unlike general FMs, which typically focus on linguistic or commonsense reasoning in structured or textual contexts, geospatial reasoning must operate over noisy, high-dimensional, and sparsely labeled spatial data. Advancing RSFMs to support such reasoning capabilities will require integrating domain knowledge, incorporating causal and temporal models, and developing architectures capable of synergistic data fusion across modalities and scales. Such models would not only perceive but also explain, predict, and reason over complex EO scenarios, enabling next-generation geospatial intelligence systems.
Explainability of RSFMs: As RSFMs are increasingly adopted in critical EO applications, enhancing their interpretability becomes essential for ensuring trust, transparency, and accountability. The black-box nature of large models poses challenges for understanding how decisions are made, particularly when integrating heterogeneous sensor modality inputs beyond optical RGB. While explainable AI (XAI) has made progress in general vision tasks, its application in RS remains limited. Developing modality-aware explainability tools tailored to RS data characteristics is a meaningful direction, enabling users to trace predictions back to geospatial and physical evidence. This will be especially crucial in sensitive applications such as environmental policy, disaster assessment, and climate monitoring.
Generalization of RSFMs: A key requirement for RSFMs is the ability to generalize across diverse spatial, temporal, and sensor domains. However, RS data often exhibits substantial distribution shifts due to seasonal changes, geographic variability, and differing acquisition conditions. Many current RSFMs still rely on fine-tuning with task-specific data, which limits their generalization capabilities. Future work should investigate domain-agnostic pretraining objectives, robust alignment strategies across modalities, and data augmentation techniques that simulate real-world variability. Improving out-of-distribution generalization will be critical for deploying RSFMs in operational settings where annotated data is scarce or unseen conditions prevail.
@@ IX Conclusion
Acknowledgments
This work was supported in part by the Council for Science, Technology and Innovation (CSTI), the Cross-ministerial Strategic Innovation Promotion Program (SIP), Development of a Resilient Smart Network System against Natural Disasters (Funding agency: NIED), JST, FOREST under Grant Number JPMJFR206S. It was also partly supported by the Ministry of Education Singapore through the Tier-2 Scheme under Project MOE-T2EP20123-0003.